OpenStudy (anonymous):

can -2x^2 + 4x +4 be done by completing the square?

hartnn (hartnn):

ANY quadratic equation can be solved by completing the square.

OpenStudy (anonymous):

how do i do that?

hartnn (hartnn):

first, try to make the co-efficient of x^2 as 1. -2x^2 + 4x +4 = -2 (x^2...........) ?

OpenStudy (anonymous):

-2(x^2 - 2x - 2)

hartnn (hartnn):

yes, thats correct. so, -2x^2 + 4x +4 =0 x^2 - 2x - 2 = 0 now, divide the co-efficient of 'x' , which is -2 , by 2 so you get -1 , if you square it, you get +1 so, x^2 - 2x - 2 = 0 ----> (x^2-2x+1) -2-1 =0 git this ?

OpenStudy (anonymous):

yeah, i got up to there but im confused what to do next

hartnn (hartnn):

sorry for late reply. now (x^2-2x+1) is a perfect square. (x- 1)^2 -3 =0 (x- 1)^2 = 3 taking square root on both sides, x-1 = +sqrt 3 or x-1 = -sqrt 3 got this ? now finding 2 values of x is pretty simple...